OpenStudy (anonymous):

A chemist is making 200 L of a solution that is 62% acid. He is mixing an 80% acid solution with a 30% acid solution. How much of the 80% acid solution will he use? @Nurali

OpenStudy (anonymous):

80% solution DATA: Amt = x liters; amount of acid is 0.8x liters ---------------------------------- 30% solution DATA: Amt = 12-x liters; amt of acid is 0.3(12-x) = 3.6-0.3x liters ---------------------------------- Amt = 12 liters; amt of acid is 0.62*12 = 7.44 liters --------------------------------------- acid + acid = acid 0.8x + 2.6-0.3x = 7.44 0.5x = 4.84 x = 9.68 liters of 80% solution in the mixture Hope I helped(: